# Improving robustness to corruptions with multiplicative weight perturbations

### v1

@@ 4.3 Benchmark results
Table 2 presents the predictive errors for the ViT-S16 / ImageNet setting, using the training setup from Beyer et al. (2022) but with only basic Inception-style preprocessing (Szegedy et al., 2016). Remarkably, DAMP can train ViT-S16 from scratch in 200 epochs to match ResNet50’s accuracy without advanced data augmentation. This is significant as ViT typically requires either extensive pretraining (Dosovitskiy et al., 2021), comprehensive data augmentation (Beyer et al., 2022), sophisticated training techniques (Chen et al., 2022), or modifications to the original architecture (Yuan et al., 2021) to perform well on ImageNet. Additionally, DAMP consistently ranks in the top 2 for corruption robustness across various test settings. Comparing Tables 1 and 2 reveals that ViT-S16 is more robust to corruptions than ResNet50 when both have similar performance on clean images.
Table 2: DAMP can train ViT-S16 on ImageNet from scratch, matching ResNet50’s accuracy in 200 epochs with only the basic Inception-style augmentations. Due to the high training cost, we report the predictive error (lower is better) of a single run for the ViT-S16 / ImageNet experiments. For each level of corruption severity in ImageNet-C and ImageNet-$\bar{\text{C}}$, we report the average error over all corruption types. All experiments use only the basic Inception-style preprocessing.
Method #Epochs ImageNet ImageNet-C ImageNet-$\bar{\text{C}}$
1 2 3 4 5 1 2 3 4 5
AdamW $100$ $28.5$ $45.4$ $56.4$ $65.8$ $76.6$ $85.1$ $48.2$ $56.9$ $64.6$ $73.8$ $79.2$
$200$ $28.7$ $46.4$ $58.0$ $67.1$ $77.4$ $85.6$ $48.6$ $57.2$ $64.9$ $74.0$ $79.4$
DAMP $100$ $25.5$ $38.9$ $48.1$ $55.9$ $67.7$ $78.7$ $41.2$ $49.0$ $56.9$ $66.3$ $71.6$
$200$ $\bf{23.7}$ $36.7$ $45.7$ $53.3$ $\bf{65.4}$ $76.8$ $39.6$ $47.6$ $55.6$ $\bf{64.9}$ $\bf{70.3}$
SAM $100$ $23.9$ $\bf{35.9}$ $45.0$ $53.2$ $65.9$ $77.8$ $39.8$ $47.9$ $55.8$ $65.3$ $70.8$
ASAM $100$ $24.0$ $\bf{35.9}$ $\bf{44.8}$ $\bf{53.0}$ $\bf{65.4}$ $\bf{76.6}$ $\bf{39.3}$ $\bf{47.3}$ $\bf{55.4}$ $65.2$ $70.9$
@@ 6 Conclusion
In this work, we demonstrate that multiplicative weight perturbations improve robustness of DNNs to a wide range of input corruptions. We introduce DAMP, a simple training algorithm that perturbs weights during training with random multiplicative noises while maintaining the same training cost as standard SGD. We further show that ASAM (Kwon et al., 2021) can be viewed as optimizing DNNs under adversarial multiplicative weight perturbations. Our experiments show that both DAMP and ASAM indeed produce models that are robust to corruptions. DAMP is also shown to improve sample efficiency of Vision Transformer, allowing it to achieve comparable performance to ResNet50 on medium size datasets such as ImageNet without extensive data augmentations. As DAMP is domain-agnostic, one future direction is to explore its effectiveness in domains other than computer vision, such as natural language processing and reinforcement learning. Another direction is to explore alternative noise distributions to the Gaussian distribution used in our work.
@@
Here we outline some limitations of this work. First, the proof of Theorem 1 assumes a simple feedforward neural network, thus it does not take into accounts modern DNN’s components such as normalization layers and attentions. Second, we did not evaluate DAMP in conjunction with advanced augmentation such as AugMix (Hendrycks et al., 2019) and AutoAugment (Cubuk et al., 2018). Finally, we only explored random Gaussian multiplicative perturbations, and there are likely more sophisticated noise distributions that could further boost corruption robustness.
@@ Appendix A Proof of Lemma 1
Here we provide the proof for Lemma 1 given the conditions and assumptions stated in Section 2:
Lemma.
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x}),\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{C_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (28)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\ell(\boldsymbol{\omega},\mathbf{g}(\mathbf{x}),y)-\ell(\boldsymbol{\omega},\mathbf{x},y)$ (29)
$\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{f}^{(h)}(\mathbf{g}(\mathbf{x}))-\mathbf{f}^{(h)}(\mathbf{x})$ (30)
We first notice that the per-sample loss $\ell(\boldsymbol{\omega},\mathbf{x},y)$ can be viewed as a function of the intermediate activation $\mathbf{f}^{(h)}(\mathbf{x})$ of layer $h$. From Assumption 2, there exists a constant $L_{h}>0$ such that:
$||\nabla_{\mathbf{f}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{g}(\mathbf{x}),y)-\nabla_{\mathbf{f}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)||_{2}\leq L_{h}||\mathbf{f}^{(h)}(\mathbf{g}(\mathbf{x}))-\mathbf{f}^{(h)}(\mathbf{x})||_{2}$ (31)
@@
$\ell(\boldsymbol{\omega},\mathbf{g}(\mathbf{x}),y)\leq\ell(\boldsymbol{\omega},\mathbf{x},y)+\left\langle\nabla_{\mathbf{f}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})\right\rangle+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})||_{2}^{2}$ (32)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq\left\langle\nabla_{\mathbf{f}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})\right\rangle+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})||_{2}^{2}$ (33)
Since the pre-activation output of layer $h+1$ is $\mathbf{z}^{(h+1)}(\mathbf{x})=\mathbf{W}^{(h+1)}\mathbf{f}^{(h)}(\mathbf{x})$, we can rewrite the inequality above as:
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x}),\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})||_{2}^{2}$ (34)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(0)}(\mathbf{x})||_{2}^{2}=||\mathbf{g}(\mathbf{x})-\mathbf{x}||_{2}^{2}\leq M$ (35)
Given that $\mathbf{f}^{(1)}(\mathbf{x})=\boldsymbol{\sigma}^{(1)}\left(\mathbf{W}^{(1)}\mathbf{f}^{(0)}(\mathbf{x})\right)$, we have:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(1)}(\mathbf{x})||_{2}^{2}=||\mathbf{f}^{(1)}(\mathbf{g}(\mathbf{x}))-\mathbf{f}^{(1)}(\mathbf{x})||_{2}^{2}\leq||\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(0)}(\mathbf{x})||_{2}^{2}$ (36)
Here we assume that the activate $\boldsymbol{\sigma}$ satisfies $||\boldsymbol{\sigma}(\mathbf{x})-\boldsymbol{\sigma}(\mathbf{y})||_{2}\leq||\mathbf{x}-\mathbf{y}||_{2}$, which is true for modern activation functions such as ReLU. Since $||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(0)}(\mathbf{x})||_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})$ such that:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(1)}(\mathbf{x})||_{2}^{2}=||\mathbf{f}^{(1)}(\mathbf{g}(\mathbf{x}))-\mathbf{f}^{(1)}(\mathbf{x})||_{2}^{2}\leq||\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(0)}(\mathbf{x})||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})}{2}||\mathbf{W}^{(1)}||_{F}^{2}$ (37)
where $||\cdot||_{F}$ denotes the Frobenius norm. Similarly, as we have proven that $||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(1)}(\mathbf{x})||_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})$ such that:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(2)}(\mathbf{x})||_{2}^{2}=||\mathbf{f}^{(2)}(\mathbf{g}(\mathbf{x}))-\mathbf{f}^{(2)}(\mathbf{x})||_{2}^{2}\leq||\mathbf{W}^{(2)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(1)}(\mathbf{x})||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})}{2}||\mathbf{W}^{(2)}||_{F}^{2}$ (38)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{f}^{(h)}(\mathbf{x})||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (39)
By combining Eqs. 34 and 39 and setting $C^{(h)}_{\mathbf{g}}(\mathbf{x})=L_{h}\hat{C}_{\mathbf{g}}^{(h)}(\mathbf{x})$, we arrive at Eq. 28. ∎
Appendix B Training with corruption
Here we present Algorithm 2 which uses corruptions as data augmentation during training, as well as the experiment results of Section 4.1 for ResNet18/CIFAR-10 and PreActResNet18/TinyImageNet settings in Figs. 4 and 5.

### v2

@@ 4.3 Benchmark results
Table 2 presents the predictive errors for the ViT-S16 / ImageNet setting, using the training setup from Beyer et al. (2022) but with only basic Inception-style preprocessing (Szegedy et al., 2016). Remarkably, DAMP can train ViT-S16 from scratch in 200 epochs to match ResNet50’s accuracy without advanced data augmentation. This is significant as ViT typically requires either extensive pretraining (Dosovitskiy et al., 2021), comprehensive data augmentation (Beyer et al., 2022), sophisticated training techniques (Chen et al., 2022), or modifications to the original architecture (Yuan et al., 2021) to perform well on ImageNet. Additionally, DAMP consistently ranks in the top 2 for corruption robustness across various test settings and has the best corruption robustness on average (last column). Comparing Tables 1 and 2 reveals that ViT-S16 is more robust to corruptions than ResNet50 when both have similar performance on clean images.
Table 3: ViT / ImageNet (IN) with MixUp and RandAugment. We train ViT-S16 and ViT-B16 on ImageNet from scratch with advanced data augmentations (DAs). We evaluate the models on IN-C, IN-$\overline{\mathrm{C}}$, IN-A, IN-D, IN-Sketch, IN-Drawing, IN-Cartoon and adversarial examples generated by FGSM. For FGSM, we use $\epsilon=2/224$. For IN-C and IN-$\overline{\mathrm{C}}$, we report the results averaged over all corruption types and severity levels. The Avg column displays the average of all previous columns except IN-Clean. These results indicate that: (i) DAMP can be combined with modern DA techniques to further enhance robustness; (ii) DAMP is capable of training large models like ViT-B16; (iii) given the same amount of training time, it is better to train a large model (ViT-B16) using DAMP than to train a smaller model (ViT-S16) using SAM/ASAM.
Model Method #Epochs Runtime Error (%) $\downarrow$
IN-Clean FGSM IN-$\overline{\mathrm{C}}$ IN-A IN-C IN-Cartoon IN-D IN-Drawing IN-Sketch Avg
ViT-S16 Dropout 500 111h 20.25 62.45 40.85 84.29 44.72 34.35 86.59 56.31 71.03 60.07
DAMP 500 111h 20.09 59.87 39.30 83.12 43.18 34.01 84.74 54.16 68.03 58.30
SAM 300 123h 20.17 59.92 40.05 83.91 44.04 34.34 85.99 55.63 70.85 59.34
ASAM 300 123h 20.38 59.38 39.44 83.64 43.41 33.82 85.41 54.43 69.13 58.58
ViT-B16 Dropout 275 123h 20.41 56.43 39.14 82.85 43.82 33.13 87.72 56.15 71.36 58.83
DAMP 275 124h 19.36 55.20 37.77 80.49 41.67 31.63 87.06 52.32 67.91 56.76
SAM 150 135h 19.84 61.85 39.09 82.69 43.53 32.95 88.38 55.33 71.22 59.38
ASAM 150 136h 19.40 58.87 37.41 82.21 41.18 30.76 88.03 51.84 69.54 57.48
ViT / ImageNet / Advanced augmentations
Table 3 presents the predictive errors of ViT-S16 and ViT-B16 on ImageNet with MixUp (Zhang et al., 2018) and RandAugment (Cubuk et al., 2020). These results indicate that DAMP can be combined with modern augmentation techniques to further improve robustness. Furthermore, using DAMP to train a larger model (ViT-B16) yields better results than using SAM/ASAM to train a smaller model (ViT-S16), given the same amount of training time.
@@ 6 Conclusion
In this work, we demonstrate that MWPs improve robustness of DNNs to a wide range of input corruptions. We introduce DAMP, a simple training algorithm that perturbs weights during training with random multiplicative noise while maintaining the same training cost as standard SGD. We further show that ASAM (Kwon et al., 2021) can be viewed as optimizing DNNs under adversarial MWPs. Our experiments show that both DAMP and ASAM indeed produce models that are robust to corruptions. DAMP is also shown to improve sample efficiency of Vision Transformer, allowing it to achieve comparable performance to ResNet50 on medium size datasets such as ImageNet without extensive data augmentations. Additionally, DAMP can be used in conjunction with modern augmentation techniques such as MixUp and RandAugment to further boost robustness. As DAMP is domain-agnostic, one future direction is to explore its effectiveness in domains other than computer vision, such as natural language processing and reinforcement learning. Another direction is to explore alternative noise distributions to the Gaussian distribution used in our work.
@@
Here we outline some limitations of this work. First, the proof of Theorem 1 assumes a simple feedforward neural network, thus it does not take into accounts modern DNN’s components such as normalization layers and attentions. Second, we only explored random Gaussian multiplicative perturbations, and there are likely more sophisticated noise distributions that could further boost corruption robustness.
@@
Acknowledgments
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI and decision no. 359567, 345604 and 341763), ELISE Networks of Excellence Centres (EU Horizon: 2020 grant agreement 951847) and UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We acknowledge the computational resources provided by Aalto Science-IT project and CSC–IT Center for Science, Finland.
@@ Appendix A Proof of Lemma 1
$\mathbf{x}^{(h)}$ :$\vcentcolon=\mathbf{f}^{(h)}(\mathbf{x})$ (19)
$\mathbf{x}_{\mathbf{g}}^{(h)}$ :$\vcentcolon=\mathbf{f}^{(h)}(\mathbf{g}(\mathbf{x}))$ (20)
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ :$\vcentcolon=\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)-\ell(\boldsymbol{\omega},\mathbf{x},y)$ (21)
$\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}$ :$\vcentcolon=\mathbf{x}^{(h)}_{\mathbf{g}}-\mathbf{x}^{(h)}$ (22)
We first notice that the per-sample loss $\ell(\boldsymbol{\omega},\mathbf{x},y)$ can be viewed as a function of the intermediate activation $\mathbf{x}^{(h)}$ of layer $h$ (see Fig. 2). From Assumption 3, there exists a constant $L_{h}>0$ such that:
$||\nabla_{\mathbf{x}^{(h)}_{\mathbf{g}}}\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)-\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)||_{2}\leq L_{h}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}$ (23)
@@
$\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)\leq\ell(\boldsymbol{\omega},\mathbf{x},y)+\left\langle\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\right\rangle+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}$ (24)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq\left\langle\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\right\rangle+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}$ (25)
Since the pre-activation output of layer $h+1$ is $\mathbf{z}^{(h+1)}(\mathbf{x})=\mathbf{W}^{(h+1)}\mathbf{f}^{(h)}(\mathbf{x})=\mathbf{W}^{(h+1)}\mathbf{x}^{(h)}$, we can rewrite the inequality above as:
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)},\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}$ (26)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}=||\mathbf{x}^{(0)}_{\mathbf{g}}-\mathbf{x}^{(0)}||_{2}^{2}=||\mathbf{g}(\mathbf{x})-\mathbf{x}||_{2}^{2}\leq M$ (27)
Given that $\mathbf{x}^{(1)}=\boldsymbol{\sigma}^{(1)}\left(\mathbf{W}^{(1)}\mathbf{x}^{(0)}\right)$, we have:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}=||\mathbf{x}^{(1)}_{\mathbf{g}}-\mathbf{x}^{(1)}||_{2}^{2}\leq||\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}$ (28)
Here we assume that the activate $\boldsymbol{\sigma}$ satisfies $||\boldsymbol{\sigma}(\mathbf{x})-\boldsymbol{\sigma}(\mathbf{y})||_{2}\leq||\mathbf{x}-\mathbf{y}||_{2}$, which is true for modern activation functions such as ReLU. Since $||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})$ such that:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}=||\mathbf{x}^{(1)}_{\mathbf{g}}-\mathbf{x}^{(1)}||_{2}^{2}\leq||\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})}{2}||\mathbf{W}^{(1)}||_{F}^{2}$ (29)
where $||\cdot||_{F}$ denotes the Frobenius norm. Similarly, as we have proven that $||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})$ such that:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(2)}||_{2}^{2}=||\mathbf{x}^{(2)}_{\mathbf{g}}-\mathbf{x}^{(2)}||_{2}^{2}\leq||\mathbf{W}^{(2)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})}{2}||\mathbf{W}^{(2)}||_{F}^{2}$ (30)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (31)
By combining Eqs. 26 and 31 and setting $C^{(h)}_{\mathbf{g}}(\mathbf{x})=L_{h}\hat{C}_{\mathbf{g}}^{(h)}(\mathbf{x})$, we arrive at Eq. 4. ∎
Appendix B Proof of Theorem 1
Proof.
From Lemma 1, we have for all $h=0,\dots,H-1$:
$\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}_{k},\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{\hat{C}_{\mathbf{g}}^{(h)}}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (33)
$\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k},\mathbf{W}^{(h)}\right\rangle}_{F}\\ +\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (34)
$\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}{\left\langle\frac{1}{N}\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k},\mathbf{W}^{(h)}\right\rangle}_{F}+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (35)
$\left[\boldsymbol{\xi}^{(h)}(\mathbf{g})\right]_{i,j}=\frac{1}{H}\frac{\left[\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k}\right]_{i,j}}{\left[\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\mathbf{x}^{(h-1)}_{k}\right]_{i,j}}$ (38)
Appendix C Training with corruption
Here we present Algorithm 2 which uses corruptions as data augmentation during training, as well as the experiment results of Section 4.1 for ResNet18/CIFAR-10 and PreActResNet18/TinyImageNet settings in Figs. 5 and 6.